OpenStudy (anonymous):

Show sequence eventually strictly increases or eventually strictly decreases. {n/ ln(n)} +infinity, n=2

OpenStudy (primeralph):

Can you draw it as is in the question? Use the "draw button".

OpenStudy (anonymous):

\[\left\{ \frac{ x }{ \ln x } \right\} ^{+\infty}_{n=2}\]

OpenStudy (anonymous):

Using n's in the equation instead of x's. Sorry. Asked to show that sequence is eventually increasing or eventually decreasing and determine smallest value of n

OpenStudy (primeralph):

Smallest value of n such that what?

OpenStudy (anonymous):

Determine smallest value of n for which the sequence is eventually strictly increasing or decreasing.

OpenStudy (primeralph):

Which is where ln(n) = 0, therefore at n = 1.

OpenStudy (anonymous):

Do you know how to calculate using to derivative rule?

OpenStudy (primeralph):

Yeah.

OpenStudy (primeralph):

Let's use x instead of n.

OpenStudy (primeralph):

Ignore my previous answer. The graph starts from 2 , so 1 is not included.

OpenStudy (anonymous):

I've gotten to \[\frac{ \ln (x) - 1 }{ (\ln (x))^{2}}\] now what do I do?

OpenStudy (primeralph):

From that, you get x = e^1 = e

OpenStudy (anonymous):

How do I find the smallest value of n from e?

OpenStudy (primeralph):

I think you should create a new question and make your prompt clearer. The smallest value is -infinity, but it looks like you're needing another type of answer. Try to create a new question and put more info. This one is already convoluted.